### v3

# Smooth band termination in 115I

Abstract
The cranked Nilsson-Strutinsky formalism is employed to investigate the band structure of 115I from low to high spin. A set of eight rotational bands with well established spins and parities are analyzed. Five bands start at low spin and are assigned to the single-particle proton orbitals ($d_{5/2}$, $g_{7/2}$), $g_{9/2}$, and $h_{11/2}$, which are close to the Z=53 Fermi surface for a deformation of $\varepsilon_{2}=0.2$. The analysis of the observed level energies relative to a rotating liquid drop and spins versus $\gamma$-ray energies allow to track the evolution of the configurations with increasing spin. It is of special interest that one band in the valence space is observed to a smooth termination at $I_{max}=67/2^{+}$.
I Introduction
The band structures of deformed nuclei with a reduced number of protons and neutrons above the $Z,N=50$ magic gaps offer optimum test benches for theoretical models conceived to describe the interplay between collective and single-particle motion. One of these models is cranked Nilsson-Strutinsky (CNS) Bengtsson and Ragnarsson (1985a), developed in particular for the interpretation of high-spin states. As pairing is neglected in this model, it has generally not been applied to low-spin states. However, the spin dependence of the difference between observed and calculated energies should mirror the pairing energy. Indeed, if it is accepted that the difference between experiment and calculations can be understood as an average pairing energy, it appears that the model can be used for all spin values, see e.g. Ragnarsson et al. (2024). An exception may be the region of the first band-crossing in prolate nuclei where the Fermi level is located at the bottom of a high-$j$ shell. However, the general features of such band-crossings are well understood from simple models, see e.g. Stephens et al. (1973) or Nilsson and Ragnarsson .
The present work is devoted to 115I, a nucleus studied in the past using $\gamma$-ray spectrometers with low and moderate efficiency Piel et al. (1985); Paul et al. (1992); Paul et al. (1994); Moon et al. (2003), but also very recently in a high-statistics experiment with the JUROGAM array Pakarinen et al. (2020), which allowed to extend the previously known bands and to identify several bands to spin values above II=30 Jodidar and Petrache . The use of clover detectors in the JUROGAM experiment was beneficial, allowing the firm assignment of spins and parities to all bands from angular correlations and polarization measurements. The spins and parities of several levels reported previously have been changed, and the level scheme has been largely revised. The experimental information on the $\gamma$-ray transitions observed in 115I will be reported in a forthcoming paper Jodidar and Petrache . In the present work we only show the bands relative to a rotating liquid drop energy in Fig. 1, and give the band-head spins and parities, the band-head energies, and the transition energies of the bands in Table 1.
Table 1: Experimental information for the bands of 115I discussed in the present work: band labels, band-head spins and parities, band-head energies (in keV), parities of the bands, and transition energies (in keV) of the in-band transitions. The new transitions, to be published in a forthcoming paper Jodidar and Petrache , are indicated wtih bold characters. For band 8, the two cascades of $E2$ transitions interconnected by dipole transitions are labeled 8e and 8o, corresponding to even and odd signature, respectively. The highest transitions of band 7 are from Ref. Paul et al. (1994). The energies of these transitions have not been observed in Ref. Jodidar and Petrache , therefore the energies are in parentheses.
1 2 3 4 5 6 7 8e 8o
$11/2^{-}$ $43/2^{-}$ $17/2^{+}$ $9/2^{+}$ $^{+} 7/2+7/2^{+}$ $43/2^{+}$ $9/2^{+}$ $11/2^{+}$
835 6479 2409 632 599 56 7096 563 876
410 1032 505 617 $\bf~654$ 736 $\bf 1028$ 633 662
516 1053 488 705 $\bf~762$ 845 $\bf 1090$ 700 732
622 $\bf 1088$ 646 781 $\bf~710$ 930 $\bf 1146$ 758 778
728 746 781 609 600 $\bf 1206$ 787 788
796 820 665 288 593 $\bf 1298$ 797 $\bf~823$
849 924 818 678 686 $\bf 1382$ $\bf~862$ $\bf~905$
893 $\bf~924$ 866 786 $\bf 1467$ $\bf~938$ $\bf~977$
1015 $\bf~736$ $\bf 1091$ 880 $\bf 1575$ $\bf 1010$ $\bf 1035$
1055 $\bf~993$ $\bf 1285$ 936 $\bf 1705$ $\bf 1071$ $\bf 1092$
1127 $\bf 1034$ 934 (1860)
1211 $\bf 1039$ $\bf~704$ (2025)
1319 $\bf 1715$ $\bf 1155$
$\bf 1382$ $\bf 1218$
$\bf 1429$ $\bf 1552$
$\bf 1652$ $\bf 2004$
$\bf 2470$
II General considerations
The band structure of 115I analyzed in the present work is composed of two negative-parity bands (bands 1 and 2) and six positive-parity bands (bands 3-8). Band 1 with negative parity and four positive-parity bands start at low spin and are interpreted as built on one-quasiparticle configurations. Band 1 built on the $11/2^{-}$ state is clearly based on the $\nu h_{11/2}$ configuration. Band 3, previously reported with negative-parity Moon et al. (2003), has in fact positive parity, as established from the polarization asymmetry of the 936-keV transition to the $27/2^{-}$ state of band 1, which indicates an $E1$ character Jodidar and Petrache . It decays via five transitions to states of the negative-parity band 1, like in the isotone 117Cs Jodidar et al. (2024), and to the positive-parity bands 4 and 6. Numerous interconnecting transitions exist between the four bands 3-6, strongly supporting similar configurations. Band 8 consisting of two nearly degenerate cascades of $E2$ transitions interconnected by $M1/E2$ dipole transitions is clearly based on a configuration with one hole in the $\pi g_{9/2}$ orbital, which give rise to strongly coupled bands, like in several nuclei of the AA=115 mass region, see e. g. Jodidar et al. (2024); Evans et al. (2006); Warring et al. (1993); Starosta et al. (2001); Zheng et al. (2021); Jodidar and Petrache (2025).
Once the configurations are assigned to the low-spin parts of the bands, one would like to understand their evolution at high spin. One can get a first idea by realizing the resemblance of the observed bands with those recently reported in the 117Cs isotone Jodidar et al. (2024), in which the low-spin bands 1, 2, 3, 4 correspond one-to-one to the bands 1, 3, 4, 8 of 115I. This is not surprising, because the two nuclei are isotones and the number of protons is different by two.
To get a first idea of the evolution of the bands with increasing spin, we analyzed the energies relative to a rotating liquid drop $E$-$E_{rld}$ as a function of spin $I$, and the spin $I$ versus the $\gamma$-ray energy $E_{\gamma}$, which are shown in Fig. 1. In the $E$-$E_{rld}$ plots one can see that all bands exhibit up-sloping patterns at low spin, which are caused by the pairing correlations. Note however, that the steep upsloping behavior is broken already at $I\approx 10$, suggesting a drastic reduction of the pairing at higher spin.
Figure 1: (a) $E$-$E_{rld}$ values for the bands of 115I. (b) Spin II versus $\gamma$-ray energy $E_{\gamma}$. Note that bands 4, 5 and 6 show a clear backbend at $E_{\gamma}\approx 0.7$ MeV, while only a minor upbend is seen for band 1.
One can distinguish three spin intervals in which bands 3-6 have regular behavior: $I\approx$0-10, $I\approx$10-20 and II>>20. For $I\approx$0-10 all curves are upsloping, with similar slope for bands 3-6 and 8, and a definitely smaller slope for band 1, which makes it yrast up to II=39/2, where band 2 becomes yrast. For $I\approx$10-20 bands 3 and 6 behave like signature partners, while bands 4 and 5 have different curvatures, but lie at excitations not much different from those of bands 3 and 6.
One can also consider that bands 4 and 5 are signature partners, as one expects the existence of four bands built on both signatures of the $(gd)$ orbitals. In fact, the energy spacing of the four [420]1/2($\alpha$$\pm$1/2) and [422]3/2($\alpha$$\pm$1/2) proton orbitals depends on the deformation and on the rotational frequency. For deformations of $\varepsilon_{2}\approx$0.2, $\gamma$$0^{{}^{\circ}}$ and frequencies of $\hbar\omega\approx$0.3 MeV, the $(gd)$ orbitals with positive signature are lowest and highest in energy, see Fig. 4, which is in qualitative agreement with the $E$-$E_{rld}$ plots, and offer an explanation of the non-observation of band 3 at low spin where it is too excited.
For II>>20, the slopes of bands 3 and 6 change drastically, which makes them rapidly yrast. A special attention should be given to band 6, which becomes rapidly non yrast at the two highest states, and behaves like a smooth terminating band Ragnarsson et al. (1995); Afanasjev et al. (1999).
The $I$ versus $E_{\gamma}$ plots in Fig. 1(b) show gradual alignments in bands 1 and 8, built on the $\pi h_{11/2}$ and $\pi g_{9/2}$ orbitals, respectively. This suggests a $\nu(h_{11/2})^{2}$ alignment in band 1, where the $\pi(h_{11/2})^{2}$ alignment is not possible because only one $h_{11/2}$ proton in present in the low-spin configurations. On the contrary, bands 3-6 built on the $(gd)$ orbitals exhibit sharp alignments at $I\approx$10 and $I\approx$22. In order to understand the nature of the observed crossings, we performed CNS calculations from which we could estimate the deformation and the excitation of the configurations close to yrast in the different spin regions, as will be presented in the following.
II.1 CNS calculations and configuration assignments
In the CNS formalism Bengtsson and Ragnarsson (1985a); Carlsson and Ragnarsson (2006), the nucleus rotates about one of its principal axes and pairing is neglected. The deformation is optimized in $(\varepsilon_{2},\gamma,\varepsilon_{4})$ for each configuration at each spin value. The configurations are labelled by the number of particles in low-$j$ and high-$j$ orbitals, in the different $\cal N$}-shells. For 115I, they can be defined relative to a 100Sn core as,
$\pi(g_{9/2})^{-p_{1}}(dg)^{p_{2}}(h_{11/2})^{p_{3}}\nu[(dg)(sd)]^{n_{1}}(h_{11/2})^{n_{2}},$
[…]
Considering only valence space configurations, after calculating all configurations with different parity and signature which become yrast in the II=0-40 spin range, we concluded that the favored configurations are $30;n_{1}n_{2}]$ and $21;n_{1}n_{2}]$ for protons, and $p_{1}p_{2};84]$, $p_{1}p_{2};93]$, and $p_{1}p_{2};A2]$ for neutrons, where A is 10 in the hexadecimal system. These four proton and six neutron configurations are then combined. The $12;n_{1}n_{2}]$ proton configurations were also analyzed, and found to be calculated high in energy, but come closest to yrast at high spin. Thus, four configurations of this type which come down in energy for II>>30 are also included. These calculated configurations are shown in Fig. 2, where they are split into positive- and negative-parity configurations to avoid a too crowded figure.
Figure 2: The calculated low-energy CNS configurations in the valence space of 110I: (a) positive and (b) negative parity. States with positive (negative) signature are drawn with filled (open) symbols.
In addition to the excitation energies, it is also interesting to follow the shape changes with increasing spin of the different configurations. To get a general overview, the deformation trajectories for all positive-parity configurations of Fig. 1(a) are shown in Fig. 3. The outcome is that almost all bands start at prolate shape ($\gamma$ within $\pm 15^{\circ}$), with $\varepsilon_{2}\geq 0.20$. With increasing spin there is a trend towards termination at $\gamma$=60, and rather different deformations in the range $\varepsilon_{2}$ $\approx$0.12-0.20.
Figure 3: Deformation trajectories for the positive-parity valence space configurations in Fig. 1(a).
A typical low-spin deformation in 115I is ($\varepsilon_{2}$=0.22, $\gamma$$0^{\circ}$), but with large fluctuations. Thus single-neutron Routhians at this deformation are drawn in Fig. 4. We can see that relative to the NN=64 shell gap, favored configurations in 115I, which has 62 neutrons, are formed with two holes in either the $h_{11/2}$ or the $(gd)$ orbitals below the gap, or with one hole in each of these groups, leading to configurations with 2, 3 or 4 neutrons in $h_{11/2}$ orbitals. For protons, the ZZ=50 core is closed for the valence space, but at a somewhat higher deformation, configurations with holes in the $g_{9/2}$ orbitals might become competitive. It follows that in the yrast region, the third proton above ZZ=50 can fill the two signatures of the $(gd)$, $h_{11/2}$ or $g_{9/2}$ orbitals.
Figure 4: (a) Single-proton and (b) single-neutron Routhians at a prolate shape typical for 115I. The orbitals are traced back to spherical shape on the left. Line convention for $(\pi,\alpha)$ orbitals: continuous for $(+,+1/2)$, dotted for (+,1/2)(+,-1/2), dashed for (,+1/2)(-,+1/2), doted-dashed for (,1/2)(-,-1/2). The occupation of the orbitals in the terminating $21_{-};9_{+}3_{-}]$ ($\pi(dg)^{2}(h_{11/2})^{1}\nu(dg)^{9}(h_{11/2})^{3}$) configuration is illustrated by filled black circles.
II.1.1 Bands 1 and 2
Band 1 is built on the $11/2^{-}$ state, suggesting that it has one proton in the lowest $h_{11/2}$ orbital with negative signature, which is largely favored relative to the positive-signature orbital, see Fig. 2(a). In CNS notation the configuration is $21_{-};84]$ at low spin, and goes through a smooth alignment at $I\approx$18, which is most naturally understood as the $\nu(h_{11/2})^{2}$ alignment. The calculated band can be followed as a smooth terminating band to II=39.5, as seen in Fig. 5, where experiment and calculations are compared. This II=39.5 state is the maximum spin of the $21_{-};84]$ configuration, where the observed band is seen up to two transitions short of termination.
Band 2 is observed for spin values II= 21.5-27.5 and is lower in energy than band 1. This might seem somewhat surprising, but gets a ready explanation from the CNS calculations as the $21_{-};A2]$ configuration, which has two less high-$j$ $h_{11/2}$ neutrons than $21_{-};84]$ and is calculated yrast above $I\approx$20, see Fig. 2(b). As noticed above, band 1 with four $h_{11/2}$ neutrons goes through a smooth crossing, in agreement with calculations, which only show a gradual variation of the slope in Fig. 5(b). In the $21_{-};A2]$ configuration assigned to band 2, aligned (or close-to-aligned) states are yrast at low energy for II=25.5, 27.5 and 29.5. The exact spin values of the energetically favored configurations depend on the relative distance between the $g_{7/2}$ and $d_{5/2}$ shells, cf. Fig. 7 below. With this in mind, it appears reasonable that band 2 is observed only to II=27.5, and not surprising that the relative energies of these states close to termination are not fully reproduced in the calculations.
[…]
Figure 6: Same as Fig. 5 but comparing of the experimental bands 3 and 6 with CNS configurations.
Figure 7: Plots of $e_{i}$ versus $m_{i}$ for protons and neutrons at $\varepsilon_{2}$=0.17, $\gamma$=60. For $Z=53$ it is shown how a favoured $I=11.5$ state is formed with one $h_{11/2}$ and two $(dg)$ protons. For NN=62, it is indicated how aligned states could be formed with 3 $h_{11/2}$ neutrons and five $(gd)$ neutron holes. Thus, a sloping Fermi surface is drawn through three $(gd)$ states. If two of these three states are empty, configurations with 5 $(gd)$ holes and three $h_{11/2}$ neurons are formed at $I_{n}$=19, 21 and 23. The red broken line shows how a maximum spin state for the same configuration but with signature $\alpha=-1/2$ is formed, i.e. for the terminating state of band 6.
Figure 8: Same as Fig. 5 but comparing of the experimental bands 7 and 8 with CNS configurations.
II.1.2 Bands 3-6
Band 6 starts at II=7/2+ at an excitation energy of 56 keV above the $5/2^{+}$ ground state. As we have seen in Fig. 1, sharp crossings occur at $I\approx$10 in bands 4, 5 and 6, while band 3 is not really observed below the crossing which might also be sharp. This is different from the smooth $\nu(h_{11/2})^{2}$ alignments observed in band 1 at $I\approx$18 and in band 8 at $I\approx$10.
Here we will concentrate on bands 3 and 6 which are observed to highest spin values. These bands are assigned to $(dg)^{3}$ proton configurations combined with neutron configurations with an even number of neutrons in the positive- and negative-parity orbitals, i.e. $(dg)^{8}(h_{11/2})^{4}$ or $(dg)^{10}(h_{11/2})^{2}$ according to the discussion above. The energy curves of these configurations are drawn in the middle panel of Fig. 6 where we note a crossing at $I\approx$15, where the $(h_{11/2})^{4}$ bands which are lower in energy at low spin, are crossed by the $(h_{11/2})^{2}$ bands which are lower at high spin. This observation suggests that the sharp crossings at II=10 are created by a change in configuration from a $(h_{11/2})^{4}$ configuration to a $(h_{11/2})^{2}$ configuration. Our interpretation is supported from the analysis of bands 1 and 2, showing that the relative energies of the $\nu(h_{11/2})^{4}$ and $\nu(h_{11/2})^{2}$ configurations are more or less reproduced in our calculations. Furthermore, both in calculations, in general, and in experiment for bands 1 and 2, the $\nu(h_{11/2})^{2}$ configuration comes relatively lower in energy with increasing spin. Thus, the combination of these two neutron configurations with the two signatures of the proton in the $(dg)$ 3/2 [422] orbital will lead to two pairs of bands which cross sharply somewhere around II=10-15. This is in general agreement with experiment, while a detailed understanding of their configurations could not really be expected in the CNS formalism. This interpretation is in line with general expectations that in high-$j$ shells, angular momentum is built more efficiently with fewer particles in the shell. This is the case for the proton $h_{11/2}$ shell in the ZZ=66-70 nuclei where with increasing spin, the number of $h_{11/2}$ protons for the yrast states decreases Bengtsson and Ragnarsson (1985b), $(h_{11/2})^{8}\rightarrow(h_{11/2})^{6}\rightarrow(h_{11/2})^{4}$ or $(h_{11/2})^{7}\rightarrow(h_{11/2})^{5}$, see also Fig. 12.11 in Ref. Nilsson and Ragnarsson .
Crossings in high-$j$ shells are generally understood as so called ’paired crossings’, i.e. a re-coupling of the spin vectors within the shell, see e. g. Stephens et al. (1973); Nilsson and Ragnarsson ; Frauendorf (1981). This is well understood in the bottom of the shell. However, as shown in Ref. Afanasjev et al. (1999), higher up in the shell when the pairing is not too strong, such crossings are naturally understood as a decrease of the number of particles in the high-$j$ shell. This interpretation challenges the more standard interpretation of crossings like that in bands 3-6 in 115I. However, it offers an explanation to the fact that only a very smooth crossing is observed in band 1, while a much sharper crossing is seen in bands 3-6.
Starting from $I\approx 10$, bands 3 and 6 look like signature partners where they are assigned to a $(dg)^{3}$ proton configuration. They go through a second crossing at $I\approx$20. One would then expect that this crossing is caused by a pair transfer, but there is no such calculated band at low energy for the highest spin values, see Fig. 2(a). Instead, all bands which are low in energy for II=20-40 have at least one proton in the $h_{11/2}$ shell. Indeed, the only reasonable interpretation of these bands which come down in energy around II=30 is the $21_{-};9_{\pm}3_{-}]$ configurations, which are clearly calculated as yrast in the spin range II=20-30, see Fig. 2.
The experimental bands 3 and 6 are compared with calculations in Fig. 6. The difference curves show a decrease with spin, which is in line with general expectations for calculations without pairing, as indicated by the smooth line drawn in Fig. 6(c). The bands are observed up to spin values just above II=30, where the calculated $21_{\pm};9_{\pm}3_{\pm}]$ bands tend to get non-collective with triaxiality sliding towards $\gamma$=60, with typical deformation $\varepsilon_{2}$=0.15-0.20.
To get a better idea about how these non-collective states are built, one can make reference to the $e_{i}$ vs. $m_{i}$ diagrams in Fig. 7, where one can see that for ZZ=53, the aligned states which are clearly favored have $I_{p}$=11.5. For neutrons, these aligned states are best characterized by neutrons in the $h_{11/2}$ shell, but with neutron holes in the $(gd)$ shells which are more than half-full. As is illustrated in Fig. 7, relatively favored aligned states can be formed for $I_{n}$=19, 21, 23. Then, combining the proton and neutron spins results in favored states at II=30.5, 32.5 and 34.5, see Fig. 7. These calculated states are consistent with the highest spin states in band 3 with low-energy II=30.5 and 32.5 sates.
Let us now consider the other signature for the [21;93] configuration corresponding to the observed band 6. It turns out that both in experiment and calculations, the energy curve is smooth up to the highest spin in the configuration, $I=35.5$. The terminating state corresponds to
$\pi(dg)^{2}_{6}(h_{11/2})^{1}_{5.5}\nu(dg)^{-5}_{10.5}(h_{11/2})^{3}_{13.5}$
where the spin values in the different groups are specified in the subscripts. Note that this is the maximum spin of signature $\alpha$=1/2-1/2 in the [21;93] configurations. To our knowledge, while many smooth terminating bands with holes in the proton core have been observed in this space Ragnarsson et al. (1995); Afanasjev et al. (1999), this is the first time that such a band within the valence space has been observed up to its maximum spin value, $I_{max}$.
II.1.3 Bands 7 and 8
Band 7 starts at II=43/2+ and is observed up to II=79/2+. As the calculated close-to-yrast configurations without holes in the $g_{9/2}$ orbital have $E$-$E_{rld}$ minima at spins much lower than II=40, see Fig. 2, the configuration of band 7 has to involve one or two holes in the $g_{9/2}$ orbital. A configuration with one hole can be excluded because it would give rise to a strongly coupled band composed of dipole and cross-over quadrupole transitions, like in band 8.
We can thus conclude that band 7 has two holes in the $\pi g_{9/2}$ orbital. The lowest energy configurations of this type are plotted in Fig. 9(a). The most favored configuration for II>>30 is $(2)3_{+}2;84]$ with signature $\alpha$=+1/2, which is different from the $\alpha$=1/2-1/2 signature of band 7. However, the $(2)3_{-}2;84]$ signature partner is not much higher in energy for spin values up to II=40 and can be assigned to band 7.
Figure 9: (a) Low-energy configurations with two proton $g_{9/2}$ holes. (b) Low-energy configurations with one proton $g_{9/2}$ hole and an even number of particles in the other groups.
Band 8 is strongly coupled and the two signatures remain degenerate up to the highest observed spin, II=23.5. The odd proton must be assigned to a high-$K$ orbital, and the only available such orbital is $g_{9/2}$, $9/2^{+}$[404]. This is consistent with the band head spin, KK=9/2. With the odd proton in the $9/2^{+}$[404] orbital, an even number of protons and neutrons should be placed in the other groups. We can thus conclude that the additional protons must be distributed as either $(dg)^{4}$ or $(dg)^{2}(h_{11/2})^{2}$ while possible neutron configurations are $(dg)^{8}(h_{11/2})^{4}$ or $(dg)^{10}(h_{11/2})^{2}$. The possible combinations of these configurations are tested in Fig. 9(b), from where one can conclude that for spin values up to $I\approx$25, [(1)40;84] bands are clearly lowest energy and should be assigned to band 8.
The bands 7 and 8 with two holes and one hole in the $g_{9/2}$ proton shell, respectively, are compared with their calculated counterpart in Fig. 8. The difference curves come out as expected with a smooth decrease with increasing spin. Indeed, if the $g_{9/2}$ orbital was lifted a few hundred keV, the difference curve would more or less coincide with the one which is used as an average for all bands of 115I, see Fig. 8(c). Note especially that while the $E$-$E_{rld}$ curves for band 8 in Fig. 6(a) and the configurations assigned to it in Fig. 6(b) do not show much resemblance, the difference curve in Fig. 6(c) comes out as expected considering that pairing is neglected in the calculations. We have also included band 1 in 113I Starosta et al. (2001). It is assigned to a configuration with two less $(dg)$ neutrons than the configuration of band 7 in 115I and leads to a very similar difference curve. This comparison illustrates the possibility to compare bands in different nuclei in the present CNS formalism Carlsson and Ragnarsson (2006) and gives some extra confidence that our assignments for bands 7 and 8 are correct.
III Summary
The level scheme of 115I has been analyzed within the CNS formalism. With ZZ=53 and NN=62, the low and intermediate spin properties are understood within the $(d_{5/2}g_{7/2})(h_{11/2})$ valence space outside the ZZ=NN=50 core. The nucleus has enough particles above closed shells to form collective bands within the valence space, but not enough neutrons to render the excitations in $(d_{3/2}s_{1/2})$ orbitals competitive in energy. The nucleus 115I exhibits a rich variety of configurations and coupling schemes when the proton particles outside the $Z=50$ are combined with neutron $h_{11/2}$ particles and $(dg)$ holes.
The bands based on the $\pi(h_{11/2})^{1}$ configuration are well understood when combined with neutron $(dg)^{8}(h_{11/2})^{4}$ and $(dg)^{10}(h_{11/2})^{2}$ configurations, leading to one smooth terminating band which is observed two transitions short of termination and another band which becomes non-collective before the $I_{max}$-value is reached.
Also the positive-parity bands with all three protons in the $(dg)$ orbitals combine with the $(dg)^{8}(h_{11/2})^{4}$ and $(dg)^{10}(h_{11/2})^{2}$ neutron configurations. In this case, the details are more difficult to understand but it is suggested that the sharp crossings in the bands at $I\approx$1010 can be understood as a crossing between $\nu(h_{11/2})^{4}$ and $\nu(h_{11/2})^{2}$ and configurations. At higher spin values above $I\approx$2020 for these positive-parity bands, the $(dg)^{3}$ proton configuration is not competitive, i.e. the $(dg)^{2}(h_{11/2})^{1}$ configuration takes over, combined with the $\nu(h_{11/2})^{3}$ configuration. In this case, the two signatures lead to one smooth terminating band and another band which becomes non-collective before the $I_{max}$-value is reached. This smooth terminating band is observed up to the maximum spin state in the configuration, i.e. to the $I_{max}$ state. While many such bands with holes in the $g_{9/2}$ orbitals have been observed for nuclei with mass numbers just below 115I Ragnarsson et al. (1995); Afanasjev et al. (1999), this is to our knowledge the first time that such a smooth terminating band has been observed to termination within the present valence space.
Going outside the valence space, the most competitive excitation is to lift protons from the $g_{9/2}$ shell of the ZZ=5050 core. With one proton excited, a strongly coupled band is formed in the $I\approx$5-25 spin range. With two protons excited, a smooth band with $E2$ transitions is observed for spin values $I\approx$20-45. These bands are well understood in their full spin range within the CNS formalism, which suggests that they evolve towards termination pretty high above yrast.